OpenStudy (anonymous):

Maggie commutes to work by train at a rate of 62 mph and then by car at a rate of 46 mph. How long will it take her to travel the 54 miles to work if she is on the train and in the car for equal periods of time? Round to the nearest tenth if necessary.

OpenStudy (anonymous):

Let x represent her travel time in hours. She is on the train and in the car for equal amounts of time. Train speed = 62 mph Car speed = 46 mph Total distance = 54 miles speed*time = distance 62x + 46x = 54 108x = 54 x = 54/108 x = .5 This would take her half an hour on the train and half an hour in the car.

OpenStudy (anonymous):

so wats the answer?

OpenStudy (anonymous):

62x + 46x = 54 108x = 54 x = 54/108 x = .5

OpenStudy (anonymous):

.5 saying that 1 is a whole hour, .5 will be half of an hour.

OpenStudy (anonymous):

oh lol Thanks

OpenStudy (anonymous):

aha no problem.